OpenStudy (anonymous):

could somebody please check my answer, log3 9x^4 - log3(3x)^2 =log3 9x^2 =2log3 9x

OpenStudy (lalaly):

\[\large{\log_3{\frac{9x^4}{9x^2}}}\]\[=\large{\log_3}{x^2}=2\log_3x\]

OpenStudy (lalaly):

you forgot to cancel out the 9

OpenStudy (anonymous):

why book gives the answer 0

OpenStudy (lalaly):

oh then the question might be (3x^2)^2?

OpenStudy (lalaly):

coz if it was (3x^2)^2 then the answer will be log3 1^2 = 2log3 1 = 0

OpenStudy (anonymous):

log3 9x4 − log3(3x)2 copied and pasted straight from the worksheet

OpenStudy (lalaly):

its either thers something wrong with the question, or the aswer given to u is wrong, the only way it can be zero is when (3x^2)^2 but in this case its 2log3 x

OpenStudy (anonymous):

thanks sensei

OpenStudy (lalaly):

your welcome:D